# A Physics-Informed Spatiotemporal Deep Learning Framework for Turbulent Systems

### v2

@@ 4 Inference
Inference is performed in two stages, wherein we first estimate the spatial parameters $\bm{W}^{\text{space}}$ and then the temporal parameters $\bm{W}^{\text{time}}$ conditional on $\widehat{\bm{W}}^{\text{space}}$. We adopt this strategy since the CAE and RNN serve two distinct (and purposely decoupled) functions. Indeed, the CAE can focus on reconstructing spatial structure, while the RNN can focus on modeling the temporal evolution in the reduced space. While joint training may offer the benefit of task-specific representations optimized end-to-end for forecasting performance, we found in practice that decoupled training yields accurate reconstructions and forecasts, while being computationally tractable. Section 4.1 describes the training and testing set, Section 4.2 describes the chosen model architecture and hyperparameters, Section 4.3 discusses the inferential process for the CAE, and Section 4.4 presents the training strategy for the temporal portion of the CRNN conditional on the estimated CAE. Inference is performed in Python 3.12.0 and Tensorflow 2.13. The PI-CRNN model was trained using 8 CPU cores for general processing and a dedicated NVIDIA A10 GPU for accelerated computation.
@@ 4.1 Training and Testing set
4.2 Hyperparameters
In the CAE, we assume that the spatial encoder and decoder each have $L=4$ layers, and the number of kernels, which are all $r_{x}^{\ell}=r_{z}^{\ell}=3$, are $K_{1}=K_{2}=K_{3}=K_{4}=64$, so the reduced space representation for any observation $\bm{Y}_{t}\in\mathbb{R}^{n_{x}\times n_{z}\times d}$ is $\tilde{\bm{Y}}_{t}\in\mathbb{R}^{16\times 16\times 64}$, $n_{x}=n_{z}=256$, resulting in a dimension reduction of 93.75%. The activation function $g$, in (3) and (5), is the leaky rectified linear unit, a modified version of the rectified linear unit that allows some negative values, in every layer except for the last, where $g$ is the hyperbolic tangent function function (Goodfellow et al. 2016). In the RNN, the context builder, $\mathcal{T}_{\text{con}}$, and sequence generator, $\mathcal{T}_{\text{gen}}$, each have $\tilde{K}=128$ kernels, where each kernel is $\tilde{r}_{x}=\tilde{r}_{z}=1$, i.e., the temporal evolution is modeled “pixel-wise” in the reduced space. For the lengths of the input and output sequences, $b$ and $a$, respectively, we leveraged the turnover time parameter $\tau_{c}$, defined as the time required for a fluid particle to complete one full convection cycle, which for this flow is $\tau_{c}=5.4s$ (or 54 observations). We selected $b=80$ and $a=60$ to ensure that both the input and output windows span at least one turnover time. This provides the model with enough temporal context to capture the dominant dynamical structures and evolve them meaningfully into the future. So the temporal model (in reduced space) takes an input sequence of length $b=80$ time steps, $\tilde{\mathcal{Y}}^{-}_{t;b}$, and outputs the next $a=60$ (reduced) observations, $\tilde{\mathcal{Y}}^{+}_{t;a}$. Further details on the formal definition of $\tau_{c}$ are available in Section S.2 of the supplementary material.
@@ 4.3 Convolutional Autoencoder Inference
The $\bm{\lambda}^{\text{space}}$ hyperparameters balance the model training and are not chosen via cross-validation. They are instead implemented via a dynamic method that automatically balances the four terms in (9) and lessens the risk that the optimization ends up at a local minimum (Wang et al. 2023). This ensures that the model does not disproportionately focus on optimizing one output at the expense of others, thereby promoting more balanced training. The approach follows established practices in multi-task learning, where adaptive loss weighting has been shown to improve convergence and final model performance (Wang et al. 2020). Technical details are available in Section S.4 of the supplementary material.
@@ 4.4 Recurrent Neural Network Physics-Informed Inference
and the vector $\bm{\lambda}^{\text{time}}=\left(\lambda^{\text{Data}}_{i},\lambda^{\text{PDE}}_{j}\right),~i\in\{u,w,p,\theta\},~j\in\{\text{Mass},~\text{Mom-}u,~\text{Mom-}w,~\text{Energy}\}$, contains weights balancing the eight terms’ respective contributions to the objective function $\mathcal{L}(\bm{W}^{\text{time}})$. The derivatives in the PDE penalty are computed using a central difference method, which we detail in the supplementary material. The dynamic loss balancing for $\bm{\lambda}^{\text{time}}$ in (10) is similar to (9) and detailed in Section S.4 of the supplementary material.
We hereto denote the forecast at time $t$ by $\widehat{\bm{Y}}_{t}=\left(\widehat{u}_{t},\widehat{w}_{t},\widehat{p}_{t},\widehat{\theta}_{t}\right)$ to simplify the notation. After training the PI-CRNN model, we predict new realizations of RBC using input sequences in the testing set $\mathcal{D}^{\text{RNN}}_{\text{test}}$ as an input and forecasting over a time horizon of $R$ time steps. The procedure to generate forecasts is outlined in Algorithm 1.
Algorithm 1 Forecasting sequences of arbitrary length
[…]
4:$t=0$
5:for $t<R$ do $\triangleright$ $R$ is the forecast horizon
6:  $\bm{\widehat{\tilde{Y}}}_{t},\bm{H}_{t},\bm{C}_{t}=\mathcal{T}_{\text{gen}}\left(\bm{\widehat{\tilde{Y}}}_{t-1},\bm{H}_{t-1},\bm{C}_{t-1}\ \middle|\ \widehat{\bm{W}}_{\text{gen}}\right)$ $\triangleright$  Reduced space prediction and context matrices
7:  $\widehat{\bm{Y}}_{t}=\mathcal{SD}\left(\bm{\widehat{\tilde{Y}}}_{t}\mid\widehat{\bm{W}}_{\text{de}}\right)$ $\triangleright$ Forecast (decompression to the original spatial scale)
8:  $t=t+1$
9:end for
[…]
5 Uncertainty Quantification
The high sensitivity of RBC with respect to the initial condition makes predicting its specific dynamics exceptionally challenging. Infinitesimally small perturbations, e.g., in thermal fluctuations or boundary layer instabilities, create diverging dynamics, thereby limiting the reliability of uncertainty quantification (UQ) over long forecasts (Leutbecher & Palmer 2008, Ahlers et al. 2009, Lorenz 1963). For this reason, we limit UQ to the first two turnover times of the forecast (corresponding to $T_{\text{uq}}=108$ time steps), beyond which the rapid divergence of trajectories renders probabilistic forecasts impractical.
We quantify the uncertainty using a bootstrap-based calibration method, which relies on conformal prediction intervals (Gammerman et al. 1998, Bonas & Castruccio 2023, Bonas et al. 2024). Conformal prediction intervals are constructed by estimating residual uncertainty using quantiles, thereby not relying on the Gaussianity assumption. Since the model was trained up to $n$, we generate an ensemble of $E=50$ forecasts over two turnover times ($T_{\text{uq}}$ total time steps) given input sequences $\mathcal{Y}^{-}_{t;b}$, $n\leq t\leq T-T_{\text{uq}}$. So, the $T_{\text{uq}}$ predicted observations are all within the testing set. Denote the $j$-th element of the long-range forecast starting at $e$ by $\widehat{\bm{Y}}_{t_{e}+j}$, $e=1,\ldots,E$. The $E$ forecasts are then grouped according to their position within the estimated sequence, $\bm{f}_{1}=\left(\widehat{\bm{Y}}_{t_{1}+1},\widehat{\bm{Y}}_{t_{2}+1},\ldots,\widehat{\bm{Y}}_{t_{E}+1}\right)$, $\bm{f}_{2}=\left(\widehat{\bm{Y}}_{t_{1}+2},\widehat{\bm{Y}}_{t_{2}+2},\ldots,\widehat{\bm{Y}}_{t_{E}+2}\right)$, etc… So $\bm{f}_{1}$ contains all the one-step ahead forecasts, $\bm{f}_{2}$ contains all the two-step ahead forecasts, and so on. Then, we compute the absolute value of the residuals for each $t$-step ahead, $\bm{r}_{t}=\left(\bm{R}_{t_{1}+t},\bm{R}_{t_{1}+t},\ldots,\bm{R}_{t_{E}+t}\right)$, where $\bm{R}_{t_{e}+t}=\left|\bm{Y}_{t_{e}+t}-\widehat{\bm{Y}}_{t_{e}+t}\right|,~\bm{R}_{t_{e}+t}=\left(\bm{R}_{u,t_{e}+t},\bm{R}_{w,t_{e}+t},\bm{R}_{p,t_{e}+t},\bm{R}_{\theta,t_{e}+t}\right),~t=1,\ldots,T_{\text{uq}},~e=1,\ldots,E$. The samples $\bm{r}_{t}$ are used to estimate the $\alpha$-quantile $\bm{q}_{j,t}$, where $j=u,w,p,\theta$, formally,
@@ 6 Results
We now present the results of our approach and first compare its spatial performance against competing statistical approaches and then the temporal one. In regard to the dimension-reduction (and subsequent reconstruction) of the data, we compare the CAE with principal component analysis (PCA), independent component analysis (ICA), fixed rank Kriging (FRK), and proper orthogonal decomposition (POD). We compare the forecast obtained using our PI-CRNN with a non-physics-informed CRNN, an echo state network (ESN), and an autoregressive integrated moving average model (ARIMA). We also assess the ability of PI-CRNN to forecast well beyond the training forecast horizon.
@@
We assess our model’s performance according to its two stated objectives: efficient reduced space representation and long-range forecasting of RBC. For spatial reconstruction, we analyze the CAE’s ability to represent high-dimensional spatial data into a reduced space and subsequently reconstruct it back onto the original dimensions. The metrics we employ are the mean squared error and the structural similarity index measure (SSIM, Wang et al. (2004)) between the reconstruction and ground truth over the testing set. The SSIM was originally conceived to quantify the similarity between images by comparing visual fidelity with respect to the original data. For long-range forecasting, we assess adherence to the governing PDE as well as physical properties of RBC that should be preserved. We focus on 1) the Nusselt number: a measure of how efficiently heat is transferred through a fluid by convection as compared to conduction; 2) the probability density functions (PDFs) of the velocity component and temperature; 3) dissipation rate of thermal fluctuations: loss of temperature variance due to heat diffusion and conduction in the fluid, respectively (Emran & Schumacher 2012). The forecast we generate is of 7 turnovers $\tau_{c}$, or 37.8 seconds, i.e., $R=378$, which is considerably longer (hence more challenging) than previous attempts to forecast similar turbulent flows (Mohan et al. 2019, Mo et al. 2019, Tang et al. 2020). A thorough explanation on the reliability of the forecast as a function of time is available in Section S.6 of the supplementary material. Technical details on each metric of assessment are also available in Section S.6 of the supplementary material. Due to RBC’s extreme sensitivity to initial conditions, even minute discrepancies, e.g., numerical perturbations, can cause forecasts to diverge rapidly over time. This inherent unpredictability makes uncertainty quantification essential. Rather than relying on point predictions alone, it is crucial to capture the range of plausible future states the system may evolve into. Therefore, in our work we also aim at assessing the quality and reliability of the predictive uncertainty provided by our PI-CRNN.
@@
The following sub-sections are organized as follows. In Section 6.1 we analyze the spatial portion of our model, while in Section 6.2 we assess the PI-CRNN forecast and its adherence to the underlying physics. Finally, in Section 6.3 we discuss the long-range forecasting ability of PI-CRNN against existing numerical methods in terms of physical accuracy and computational cost.
@@ 6.2 Temporal Analysis
We compare our forecast of $\tau_{c}=7$ turnover times with the ESN, CRNN, and ARIMA. The CRNN model is identical to PI-CRNN, but is not informed with the governing PDE, so the loss function we minimize is similar to (10) but without the “PDE” terms. The ESN is a type of recurrent neural network with a similar structure as a LSTM network, but with considerably less computational demand dictated by the sparseness of its parameter matrices (McDermott & Wikle 2019). The main advantage of ESNs is the absence of backpropagation through time (which is the computational bottleneck of LSTM and ConvLSTM), since only the output weights are trained. Since we perform physics-informed inference, we hereby denote it by “PI-ESN”. The forecast is generated by recursively applying the ESN $R$ times. Lastly, the ARIMA model is one of the most popular approaches in time series analysis thanks to its simplicity and ease of implementation. Similar to the ESN, we generate a forecast of length $R$ by recursively applying ARIMA. Technical details on PI-ESN, CRNN, and ARIMA are available in Section S.6.2 of the supplementary material.
@@
Figure 3 presents a sequence of snapshots from the temperature forecast $\widehat{\theta}_{t}$ over the course of the first two turnover times, corresponding to approximately 11 seconds of physical time. Each panel shows the predicted temperature field at different times, highlighting the evolution of the thermal structures. Two fully developed plumes can be clearly observed rising from the bottom boundary and interacting with the surrounding flow. PI-CRNN is able to capture the key dynamical features of plume formation, growth, and lateral motion over time. Although the mean squared error with respect to the true data increases gradually with forecast lead time, indicating some accumulation of small deviations, the residuals of the governing equations remain relatively constant, especially over the first two turnover times (see Figure S10 of the supplementary material). This suggests that while the forecast slowly diverges from the realized RBC for $t>n$, the model consistently preserves the underlying physical constraints. An animation showing the full temporal evolution is available in the online code associated with the manuscript.
Figure 3: Snapshots of the temperature forecast over the first turnover time ($\approx$ 5.5s), where the spatiotemporal dynamics remain strongly dependent on the input sequence. An animation is available in the online code associated with the manuscript.
@@ 7 Conclusion
The code for this work can be found in the following GitHub repository: github.com/papersblindreview/PI_CRNN_RBC_2025. The numerical solver for the RBC used in this work can be freely downloaded from git.uwaterloo.ca/SPINS/SPINS_main.

### v3

@@ 4 Inference
Inference is performed in two stages, wherein we first estimate the spatial parameters $\bm{W}^{\text{space}}$ and then the temporal parameters $\bm{W}^{\text{time}}$ conditional on $\widehat{\bm{W}}^{\text{space}}$. We adopt this strategy since the CAE and RNN serve two distinct (and purposely decoupled) functions. Indeed, the CAE can focus on reconstructing spatial structure, while the RNN can focus on modeling the temporal evolution in the reduced space. While joint training may offer the benefit of task-specific representations optimized end-to-end for forecasting performance, we found in practice that decoupled training yields accurate reconstructions and forecasts, while being computationally tractable. Section 4.1 describes the training and testing set, Section 4.2 discusses the inferential process for the CAE, and Section 4.3 presents the training strategy for the temporal portion of the CRNN conditional on the estimated CAE. Inference is performed in Python 3.12.0 and Tensorflow 2.13. The model was trained using 8 CPU cores for general processing and a dedicated NVIDIA A10 GPU for accelerated computation.
@@ 4.2 Convolutional Autoencoder Inference
The $\bm{\lambda}^{\text{space}}$ hyperparameters balance the model training and are not chosen via cross-validation. They are instead implemented via a dynamic method that automatically balances the four terms in (9) and lessens the risk that the optimization ends up at a local minimum (Wang et al. 2023a). This ensures that the model does not disproportionately focus on optimizing one output at the expense of others, thereby promoting more balanced training. The approach follows established practices in multi-task learning, where adaptive loss weighting has been shown to improve convergence and final model performance (Wang et al. 2020). Technical details are available in Section S.4 of the supplementary material.
@@ 4.3 Recurrent Neural Network Physics-Informed Inference
and the vector $\bm{\lambda}^{\text{time}}=\left(\lambda^{\text{Data}}_{i},\lambda^{\text{PDE}}_{j}\right),~i\in\{u,w,p,\theta\},~j\in\{\text{Mass},~\text{Mom-}u,~\text{Mom-}w,~\text{Energy}\}$, contains weights balancing the eight terms’ respective contributions to the objective function $\mathcal{L}(\bm{W}^{\text{time}})$. The derivatives in the PDE penalty are computed using a central difference method, detailed in the supplementary material. The dynamic loss balancing for $\bm{\lambda}^{\text{time}}$ in (10) is similar to (9) and detailed in Section S.4 of the supplementary material.
We hereto denote the forecast at time $t$ by $\widehat{\bm{Y}}_{t}=\left(\widehat{u}_{t},\widehat{w}_{t},\widehat{p}_{t},\widehat{\theta}_{t}\right)$ to simplify the notation. After training the PI-CRNN model, we predict new realizations of RBC using input sequences in the testing set $\mathcal{D}^{\text{RNN}}_{\text{test}}$ as an input and forecasting over a time horizon of $R$ time steps. The procedure to generate forecasts is outlined in Algorithm S3 of the supplementary material.
@@ 5 Results
We apply the PI-CRNN model described in Section 3 to the dataset in Section 2 and address questions Q1 and Q2 in the same Section. We fist focus on Q1 and compare our model’s spatial performance against competing statistical approaches and then the temporal one. In regard to the dimension-reduction (and subsequent reconstruction) of the data, we compare the CAE with principal component analysis (PCA; Jolliffe (2002)), independent component analysis (ICA; Comon (1994)), fixed rank Kriging (FRK; Cressie and Johannesson (2008)), and proper orthogonal decomposition (POD; Berkooz et al. (1993)). Details on PCA, ICA, FRK and POD are available in Section S.6.1 of the supplementary material. Then, we address Q2 by comparing the forecast obtained using our PI-CRNN with a non-physics-informed CRNN, an echo state network (ESN), and an autoregressive integrated moving average model (ARIMA). We also assess the ability of PI-CRNN to forecast well beyond the training forecast horizon.
@@
We assess our model’s performance according to its two stated objectives: efficient reduced space representation and long-range forecasting of RBC. For spatial reconstruction, we analyze the CAE’s ability to represent high-dimensional spatial data into a reduced space and subsequently reconstruct it back onto the original dimensions. The metrics we employ are the mean squared error and the structural similarity index measure (SSIM, Wang et al. (2004)) between the reconstruction and ground truth over the testing set. The SSIM was originally conceived to quantify the similarity between images by comparing visual fidelity with respect to the original data. For long-range forecasting, we assess adherence to the governing PDE as well as physical properties of RBC that should be preserved. We focus on 1) the Nusselt number: a measure of how efficiently heat is transferred through a fluid by convection as compared to conduction; 2) the probability density functions (PDFs) of the velocity component and temperature; 3) dissipation rate of thermal fluctuations: loss of temperature variance due to heat diffusion and conduction in the fluid, respectively (Emran and Schumacher 2012). The forecast we generate is of 7 turnovers $\tau_{c}$, or 37.8 seconds, i.e., $R=378$, which is considerably longer (hence more challenging) than previous attempts to forecast similar turbulent flows (Mohan et al. 2019, Mo et al. 2019, Tang et al. 2020). A thorough explanation on the reliability of the forecast as a function of time is available in Section S.6 of the supplementary material. Technical details on each metric of assessment are also available in Section S.6 of the supplementary material. Due to RBC’s extreme sensitivity to initial conditions, even minute discrepancies, e.g., numerical perturbations, can cause forecasts to diverge rapidly over time. This inherent unpredictability makes uncertainty quantification essential. Rather than relying on point predictions alone, it is crucial to capture the range of plausible future states the system may evolve into. Therefore, in our work we also aim at assessing the quality and reliability of the predictive uncertainty provided by our PI-CRNN.
@@
The following sub-sections are organized as follows. In Section 5.1 we analyze the spatial portion of our model, while in Section 5.2 we assess the PI-CRNN forecast and its adherence to the underlying physics. Finally, in Section 5.3 we discuss the long-range forecasting ability of PI-CRNN against existing numerical methods and in Section 5.4 we discuss limitations.
@@ 5.2 Temporal Analysis
We compare our forecast of $\tau_{c}=7$ turnover times with the ESN, CRNN, and ARIMA. The CRNN model is identical to PI-CRNN, but is not informed with the governing PDE, so the loss function we minimize is similar to (10) but without the “PDE” terms. The ESN is a type of recurrent neural network with a similar structure as a LSTM network, but with considerably less computational demand dictated by the sparseness of its parameter matrices (McDermott and Wikle 2019). The main advantage of ESNs is the absence of backpropagation through time (which is the computational bottleneck of LSTM and ConvLSTM), since only the output weights are trained. Since we perform physics-informed inference, we hereby denote it by “PI-ESN”. The forecast is generated by recursively applying the ESN $R$ times. Lastly, the ARIMA model is one of the most popular approaches in time series analysis thanks to its simplicity and ease of implementation. Similar to the ESN, we generate a forecast of length $R$ by recursively applying ARIMA. Technical details on PI-ESN, CRNN, and ARIMA are available in Section S.6.2 of the supplementary material.
@@
Figure 3 presents a sequence of snapshots from the temperature forecast $\widehat{\theta}_{t}$ at selected times. The first three snapshots correspond to predictions within the first turnover time, highlighting the model’s short-term accuracy, while the last two panels show long-term predictions, where errors with respect to the DNS data have accumulated. PI-CRNN is able to capture the key dynamical features of plume formation, growth, and lateral motion over time. Although the mean squared error with respect to the true data increases gradually with forecast lead time, indicating some accumulation of small deviations, the residuals of the governing equations remain relatively constant, especially over the first two turnover times (Figures S8-S9 of the supplementary material). This suggests that while the forecast slowly diverges from the realized RBC for $t>n$, the model consistently preserves the underlying physical constraints. An animation showing the full temporal evolution is available in the online code associated with the manuscript.
Figure 3: Snapshots of the temperature forecast at selected times. The first three panels correspond to predictions within the first turnover time, highlighting the model’s short-term accuracy, while the last two panels show long-term predictions, where errors have accumulated and deviations from the DNS trajectory become more apparent. An animation is available in the online code associated with the manuscript.
@@ 5.3 PI-CRNN as a Surrogate
5.4 Limitations
PI-CRNN presents a few structural shortcomings that could be mitigated in future research. First and foremost, maintaining physical consistency over longer horizons remains a key limitation: small violations of the governing equations in the learned context dynamics can accumulate, leading to drift from physically plausible states in the strongly chaotic regime. Extending PI-CRNN’s reliability horizon would be beneficial for applications in which the relevant physics unfolds over longer timescales, e.g., particle dispersion. The approach is also computationally expensive to train, the result of a large parameter space and the physics-based loss terms that require evaluating spatial derivatives on fine grids, resulting in substantial memory and time costs. Finally, the model has been trained and validated for only a single set of physical parameters (one combination of Rayleigh and Prandtl numbers), which restricts its ability to generalize across different convection regimes. Incorporating multiple regimes would enable systematic analysis of particle transport, dispersion, and accumulation under varying flow conditions, allowing robust comparison of transport statistics that are relevant to real-world geophysical and industrial convection systems.
@@ 6 Conclusion
The code for this work can be found in the following GitHub repository: github.com/Env-an-Stat-group/26.Menicali.JASA. The numerical solver for the RBC used in this work can be freely downloaded from git.uwaterloo.ca/SPINS/SPINS_main.